Find the slope from the equation.
-30+10y=-2x

Answer by rapaljer(4671)

Solve for y:

-30+10y= -2x

Add +30 to each side:
10y = -2x+30

Divide both sides by 10:
y= -2/10x +30/10
y=-1/5x +3

The slope is the coefficient of x, which is -1/5.

R^2

Dr. Robert J. Rapalje